In Drosophila, red eyes is a wild-type phenotype. Several different genes affect eye color. One allele causes purple eyes, and a different allele causes sepia eyes. Both of these are recessive to red eye color. When flies with purple eyes were crossed to flies with sepia eyes, all F1 flies had red eyes. When the F1 offspring were allowed to mate, the F2 flies were as follows:

162 purple eyes

172 sepia eyes

58 purplish sepia eyes

478 red eyes

Explain this pattern of inheritance. Use a chi square test to test the goodness of fit of your hypothesis with the data
Answer #1

Answer:

Dihybrid F2 phenotypic ratio = 9:3:3:1

Dominant/dominant = 9/16

Dominant/recessive = 3/16

Recessive / dominant = 3/16

Recessive / recessive = 1/16

Total progeny = 870

purple eyes = 870 * 3/16 = 163.125

sepia eyes = 870 * 3/16 = 163.125

purplish sepia eyes= 870 * 1/16 = 54.375

red eyes = 870 * 9/16 = 489.375

Phenotype

Observed(O)

Expected (E)

O-E

(O-E)2

(O-E)2/E

Red eyes

478

489.38

-11.375

129.391

0.264

purple eyes

162

163.13

-1.125

1.266

0.008

sepia eyes

172

163.13

8.875

78.766

0.483

purplish sepia eyes

58

54.38

3.625

13.141

0.242

Total

870

870

0.997

Chi-square value = 0.997

Degrees of freedom = Number of phenotypes – 1

Df = 4-1=3

Critical value 7.81

The chi-square value of 0.997 is less than the critical value of 7.81. We can the accept null hypothesis and the data fit into the hypothesis.